# Multitype entanglement dynamics induced by exceptional points

### v1

@@ Abstract
As a most important feature of non-Hermitian systems, exceptional points (EPs) lead to a variety of unconventional phenomena and applications. Here we discover that multitype entanglement dynamics can be induced by engineering different orders of EP. By studying a generic model composed of two coupled non-Hermitian qubits, we find that diverse entanglement dynamics on the two sides of the fourth-order EP (EP4) and second-order EP (EP2) can be observed simultaneously in the weak coupling regime. With the increase of the coupling strength, the EP4 is replaced by an additional EP2, leading to the disappearance of the entanglement dynamics transition induced by EP4 in the strong coupling regime. Considering the case of Ising type interaction, we also realize EP-induced entanglement dynamics transition without the driving field. Our study paves the way for the investigation of EP-induced quantum effects and applications of EP-related quantum technologies.
@@ I Introduction
Non-Hermitian (NH) Hamiltonians provide an effective method to describe physical systems that exchange energy, information or particles with the environment Bender and Boettcher (1998); Heiss (2004). For instance, NH Hamiltonians have been applied to describe optical Rüter et al. (2010); El-Ganainy et al. (2018); Hodaei et al. (2017), acoustic Zhu et al. (2018); Liu et al. (2018) and magnetic systems Yang et al. (2020); Wang et al. (2019) with gain and loss or asymmetric couplings. These non-Hermitian systems have unprecedented features induced by non-HermiticityBender and Boettcher (1998). As a most peculiar example, exceptional points (EPs) describe the coalescence of the eigenstates and the degeneracy of the eigenvalues, indicating the phase transition from the parity-time ($\mathcal{PT}$) symmetric phase to the $\mathcal{PT}$-symmetry-broken phase Heiss (2004); Bergholtz et al. (2021); Özdemir et al. (2019); Parto et al. (2021); Miri and Alù (2019). EPs play an important role in many remarkable physical phenomena and functional applications. For example, the nonlinear perturbation response of EPs induces a variety of unconventional effects, including nonreciprocal light propagation Peng et al. (2014a); Chang et al. (2014), loss-induced transparency Peng et al. (2014b); Zhang et al. (2018) and coherent perfect absorption Sun et al. (2014); Wang et al. (2021), making them promising platforms for wireless power transfer Assawaworrarit and Fan (2020); Assawaworrarit et al. (2017), single-mode lasers Feng et al. (2014) and so on.
@@
In addition to investigating EPs in classical systems, the extension of EPs to quantum regime has attracted much attention. Based on the methods of extended Hilbert space Wu et al. (2019) and quantum trajectories Naghiloo et al. (2019) to construct effective NH Hamiltonians, EPs have been achieved in various quantum platforms such as dissipative photon systems Xiao et al. (2020); Öztürk et al. (2021), superconducting circuits Naghiloo et al. (2019); Chen et al. (2021); McDonald and Clerk (2020); Chen et al. (2022), single ion traps Ding et al. (2021), thermal atom ensembles Cao et al. (2020); Liang et al. (2023), cold atoms Li et al. (2019); Lee and Chan (2014) and nitrogen-vacancy color-centers Wu et al. (2019); Pick et al. (2019). Quantum effects including photon blockadeHuang et al. (2022); Li et al. (2023a); Zuo et al. (2022); Yuan et al. (2023) and topological quantum state controlLiu et al. (2021); Gong et al. (2018); Xu et al. (2016) have been achieved by engineering the EPs. As a most important quantum property, EP-induced exceptional entanglement behaviors have also been observed in very recent works. For example, the occurrence of the entanglement transition at the EP has been observed in a NH system composed of qubit coupled with photonsHan et al. (2023). Entanglement generation between two NH qubits can be accelerated by approaching the EP of the systemLi et al. (2023b). Replacing one NH qubit with unitary qubit, entanglement can be maximized at the EPKumar et al. (2022). However, since most work have only discussed the case of second-order EPs, the relationship between the orders of EPs and entanglement remains unknown.
@@
Here, we show that multitype entanglement dynamics can be realized by engineering different orders of EPs, by analyzing a generic model composed of two coupled NH qubits. The inter-qubit coupling will lower the order of EP from the original fourth-order EP (EP4) to a second-order EP (EP2). However, we find that disparate entanglement dynamics on the two sides of the original EP4 and EP2 can be observed simultaneously in the case of weak coupling (i.e., the coupling strength is much smaller than the dissipation rate), indicating different types of entanglement behaviors can be induced by tuning the order of EP. When enhancing the coupling strength to the strong coupling regime (i.e., the coupling strength and dissipation rate are on the same order), the original EP4 is replaced by an additional EP2, leading to the disappearance of the EP4-engineered entanglement transition. Taking Ising type interaction as an example, we also find EP-induced entanglement transition in the absence of the driving field. Our scheme is universal for a variety of quantum systems, such as superconducting circuits Blais et al. (2020); Xiang et al. (2013); Devoret and Schoelkopf (2013) and ion traps Monroe and Kim (2013); Pino et al. (2021). This study paves the way for the investigation of quantum effects induced by EPs, and provides opportunities for designing EP-based quantum devices.
@@
Figure 1: Schematic diagram of the multitype entanglement dynamics (I, II, III) induced by EPs with different orders. The regions of PTB and PTS represent the $\mathcal{P}\mathcal{T}$-symmetry-broken phase and the $\mathcal{P}\mathcal{T}$-symmetric phase, respectively. (a) The system exhibits a fourth-order EP (EP4) without coupling and no entanglement will be generated in this case. (b) In the weak coupling regime (the coupling strength $\xi$ is much smaller than the dissipation rate $\gamma$ of the qubits, i.e., $\xi\ll\gamma$), the system exhibits the original EP4 ($\text{EP}4^{\prime}$) and an existing second-order EP (EP2). (c) In the strong coupling regime (i.e, $\xi\sim\gamma$), the original EP4 disappears and an additional EP2 will emerge. The definition of type I, II and III denote three different types of entanglement dynamics. The entanglement behavior of type I is monotonically increasing and eventually reaching a stable value, type II corresponds to the entanglement behavior of continuous oscillations over the entire time domain, and type III combines the characteristics of the above two different entanglement behaviors.
@@ II Model
$H=\sum\limits_{j=1,2}{\left({{\delta}_{j}}-\frac{i{{\gamma}_{j}}}{2}\right)\sigma_{j}^{-}}\sigma_{j}^{+}+{{\Omega}_{j}}\sigma_{j}^{x}+{{H}_{\operatorname{int}}},$ (1)
@@
where ${{\delta}_{j}}$ represents the frequency detuning of the driving field from the qubit transition frequency, ${{\gamma}_{j}}$ and ${{\Omega}_{j}}$ is the energy decay rate of ${{\left|b\right\rangle}_{j}}$ and driving amplitude, respectively. The Pauli operators in terms of the quantum energy levels-with no classical analogs- ${{\left|b\right\rangle}_{j}}$ and ${{\left|a\right\rangle}_{j}}$ as $\sigma_{j}^{+}={{\left|a\right\rangle}_{j}}\left\langle b\right|,\sigma_{j}^{-}={{\left|b\right\rangle}_{j}}\left\langle a\right|$ and $\sigma_{j}^{x}={{\left|b\right\rangle}_{j}}\left\langle a\right|+{{\left|a\right\rangle}_{j}}\left\langle b\right|,j\in\{1,2\}$. In the following analysis, we take ${{\gamma}_{1}}={{\gamma}_{2}}=\gamma$ and ${{\Omega}_{1}}={{\Omega}_{2}}=\Omega$ for simplicity and assumed that ${{\delta}_{j}=0}$ (i.e., the drives are resonant with two qubits). The third term represents the interaction between the two NH qubits. When the two qubits are coupled via dipolar interaction, the Hamiltonian can be described as ${{H}_{\operatorname{int}}}=g\left(\sigma_{1}^{\dagger}{{\sigma}_{2}}+{{\sigma}_{1}}\sigma_{2}^{\dagger}\right)$, which can be realized in various systems such as superconducting circuit systems Naghiloo et al. (2019); Dalmonte et al. (2015) and Rydberg atom systems de Léséleuc et al. (2019). In addition to this interaction type, there is another type of interaction between the two qubits or spins, i.e, Ising type interaction, which can be represented as ${{H}_{\operatorname{int}}}=-\xi{{\left(\sigma_{x}^{1}+\sigma_{x}^{2}\right)}^{2}}$. Here $g$ and $\xi$ denote the coupling strength, respectively. Ising type interaction has been widely studied in hybrid spin-mechanical system Pan et al. (2023); Li et al. (2016), the quantum spin lattice systems Dür et al. (2005); Britton et al. (2012) and hybrid circuit systemsBohnet et al. (2016); Block et al. (2022).
III EPs with different orders
@@
In the following we mainly focus on the case of Ising type interaction (the calculation of the dipolar interaction case can be found in Sec. VI). Due to ${{\delta}_{j}=0}$, the system Hamiltonian in Eq. (1) reduces to a so-called passive $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian Guo et al. (2009) can be written as $H={{H}_{\mathcal{PT}}}-\left(i\gamma/2\right)I$ with $I$ being the unitary matrix. The Hamiltonian ${{H}_{\mathcal{PT}}}$ can be written as
@@
$\begin{split}{{H}_{\mathcal{PT}}}=&\left(\begin{matrix}J+\frac{i\gamma}{2}&\Omega&\Omega&J\\ \Omega&J&J&\Omega\\ \Omega&J&J&\Omega\\ J&\Omega&\Omega&J-\frac{i\gamma}{2}\\ \end{matrix}\right),\end{split}$ (2)
where $J=-2\xi$. $H_{\mathcal{PT}}$ satisfies $\mathcal{P}\mathcal{T}$ symmetry, where the parity operator is given as $\mathcal{P}=\sigma_{1}^{x}\sigma_{2}^{x}$.
@@
The eigenvalues of the system without coupling, i.e., $\xi=0$ ($g=0$), are ${{\lambda}_{0,1}}=-i\gamma/2$ and ${{\lambda}_{\pm}}=-i\gamma/2\pm i\sqrt{16{{\Omega}^{2}}-{{\gamma}^{2}}}/2$, indicating the system exists a four-order EP (EP4) if $\Omega={{\Omega}_{\text{EP4}}}=\gamma/4$ [Fig. 2(a) and  2(d)]. The region of $\mathcal{P}\mathcal{T}$-symmetric phase (PTS) ($\Omega>{{\Omega}_{\text{EP4}}}$) is on the right-hand side of the EP4 while the region of $\mathcal{P}\mathcal{T}$-symmetry-broken phase (PTB) ($\Omega<{{\Omega}_{\text{EP4}}}$) is on the left-hand side of the EP4 [Fig. 1(a)]. The order of EP will be lowered from 4 to 2 when adding the qubit-qubit coupling. As shown in Fig. 2(b) and  2(e), an EP2 can be found at ${{\Omega}_{\text{EP2}}}/\gamma=0.283$ when choosing $\xi/\gamma=0.006$ as an example. In the weak coupling regime (i.e., $\xi\ll\gamma$), the coupling can also be considered as a perturbation. The original EP4 (${{\Omega}_{\text{EP}4^{\prime}}}={{\Omega}_{\text{EP}4}}$) and existing EP2 will form a new phase (${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$)[Fig. 1(b)], as different entanglement dynamics can be observed in this region, which will be mainly discussed in the following section. Interestingly, an additional EP2 will appear in the system when enhancing the coupling strength to the strong coupling regime (i.e., $\xi\sim\gamma$). As demonstrated in Fig. 2(c) and  2(f), one EP2 appears at the critical value $\Omega_{\text{EP2}}^{(1)}/\gamma=0.626$ and an additional EP2 is at $\Omega_{\text{EP2}}^{(2)}/\gamma=1.34$ when choosing $\xi/\gamma=0.5$. The region of PTB occurs at $\Omega_{\text{EP2}}^{(1)}<\Omega<\Omega_{\text{EP2}}^{(2)}$ and the regions of PTS take place at $\Omega<\Omega_{\text{EP2}}^{(1)}$ and $\Omega>\Omega_{\text{EP2}}^{(2)}$ [Fig. 1(c)].
@@
Figure 2: Real (a, b, c) and imaginary (d, e, f) parts of eigenvalues for different Ising type interaction strength. (a) Real and (d) imaginary parts of the eigenvalues for $\xi/\gamma=0$. And examples for weak coupling regime for $\xi/\gamma=0.006$ (b, e) and strong coupling regime for $\xi/\gamma=0.5$ (c, f), respectively. The two qubits are assumed to have the same drive amplitude $\Omega$ and the same decay rates $\gamma$. The orange and white regions represent the PTB and PTS, respectively. The new phase is formed by the original EP4 and existing EP2 (i.e., ${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$) in the case of weak coupling is illustrated in the blue region.

### v2

@@ Abstract
As a most important feature of non-Hermitian systems, exceptional points (EPs) lead to a variety of unconventional phenomena and applications. Here, we study a generic model composed of two coupled non-Hermitian qubits, the EPs can be easily obtained in this system by adjusting the driving amplitude applied to the qubits. The diverse entanglement dynamics on the two sides of the original fourth-order EP (EP4’) and second order EP (EP2) can be observed simultaneously in the weak coupling regime. With the increase of the coupling strength, the EP4’ is replaced by an additional EP2, leading to the disappearance of the entanglement dynamics changing induced by EP4’ in the strong coupling regime. Considering the case of Ising type interaction, we also realize EP-induced entanglement dynamics changing without the driving field. Our study paves the way for the investigation of EP-induced quantum effects and applications of EP-related quantum technologies.
@@ I Introduction
Non-Hermitian (NH) Hamiltonians provide an effective method to describe physical systems that exchange energy, information or particles with the environment Bender and Boettcher (1998); Heiss (2004). For instance, NH Hamiltonians have been applied to describe optical C. E. Rüter, K. G. Makris, R. El-Ganainy, D. N. Christodoulides, M. Segev, and D. Kip (2010); R. El-Ganainy, K. G. Makris, M. Khajavikhan, Z. H. Musslimani, S. Rotter, and D. N. Christodoulides (2018); H. Hodaei, A. U. Hassan, S. Wittek, H. Garcia-Gracia, R. El-Ganainy, D. N. Christodoulides, and M. Khajavikhan (2017), acoustic W. Zhu, X. Fang, D. Li, Y. Sun, Y. Li, Y. Jing, and H. Chen (2018); T. Liu, X. Zhu, F. Chen, S. Liang, and J. Zhu (2018) and magnetic systems Y. Yang, Y.-P. Wang, J. W. Rao, Y. S. Gui, B. M. Yao, W. Lu, and C. M. Hu (2020); Y.-P. Wang, J. W. Rao, Y. Yang, P.-C. Xu, Y. S. Gui, B. M. Yao, J. Q. You, and C. M. Hu (2019) with gain and loss or asymmetric couplings. These non-Hermitian systems have unprecedented features induced by non-HermiticityBender and Boettcher (1998). As a most peculiar example, exceptional points (EPs) describe the coalescence of the eigenstates and the degeneracy of the eigenvalues, indicating the phase transition from the parity-time ($\mathcal{PT}$) symmetric phase to the $\mathcal{PT}$-symmetry-broken phase Heiss (2004); E. J. Bergholtz, J. C. Budich, and F. K. Kunst (2021); S. K. Özdemir, S. Rotter, F. Nori, and L. Yang (2019); M. Parto, Y. Liu, B. Bahari, M. Khajavikhan, and D. N. Christodoulides (2021); M. A. Miri and A. Alù (2019). EPs play an important role in many remarkable physical phenomena and functional applications. For example, the nonlinear perturbation response of EPs induces a variety of unconventional effects, including nonreciprocal light propagation et al. (2014); L. Chang, X. S. Jiang, S. Y. Hua, C. Yang, J. M. Wen, L. Jiang, G. Y. Li, G. Z. Wang, and M. Xiao (2014), loss-induced transparency B. Peng, S. K. Özdemir, S. Rotter, H. Yilmaz, M. Liertzer, F. Monifi, C. M. Bender, F. Nori, and L. Yang (2014); H. Zhang, F. Saif, Y. Jiao, and H. Jing (2018) and coherent perfect absorption Y. Sun, W. Tan, H.-q. Li, J. Li, and H. Chen (2014); C. Q. Wang, W. R. Sweeney, A. D. Stone, and L. Yang (2021), making them promising platforms for wireless power transfer S. Assawaworrarit and S. H. Fan (2020); S. Assawaworrarit, X. F. Yu, and S. H. Fan (2017), single-mode lasers L. Feng, Z. J. Wong, R. M. Ma, Y. Wang, and X. Zhang (2014) and so on.
@@
In addition to investigating EPs in classical systems, the extension of EPs to quantum regime has attracted much attention. Based on the methods of extended Hilbert space Y. Wu, W. Q. Liu, J. P. Geng, X. R. Song, X. Y. Ye, C. K. Duan, X. Rong, and J. F. Du (2019) and quantum trajectories M. Naghiloo, M. Abbasi, Y. N. Joglekar, and K. W. Murch (2019) to construct effective NH Hamiltonians, EPs have been achieved in various quantum platforms such as dissipative photon systems L. Xiao, T. S. Deng, K. K. Wang, G. Y. Zhu, Z. Wang, W. Yi, and P. Xue (2020); F. E. Öztürk, T. Lappe, G. Hellmann, J. Schmitt, J. Klaers, F. Vewinger, J. Kroha, and M. Weitz (2021), superconducting circuits M. Naghiloo, M. Abbasi, Y. N. Joglekar, and K. W. Murch (2019); W. J. Chen, M. Abbasi, Y. N. Joglekar, and K. W. Murch (2021); McDonald and Clerk (2020); W. Chen, M. Abbasi, B. Ha, S. Erdamar, Y. N. Joglekar, and K. W. Murch (2022), single ion traps L. Y. Ding, K. Y. Shi, Q. X. Zhang, D. N. Shen, X. Zhang, and W. Zhang (2021), thermal atom ensembles W. Cao, X. Lu, X. Meng, J. Sun, H. Shen, and Y. Xiao (2020); C. Liang, Y. Tang, A.-N. Xu, and Y.-C. Liu (2023), cold atoms J. M. Li, A. K. Harterg, J. Liu, L. de Melo, Y. N. Joglekar, and L. Luo (2019); Lee and Chan (2014) and nitrogen-vacancy color-centers Y. Wu, W. Q. Liu, J. P. Geng, X. R. Song, X. Y. Ye, C. K. Duan, X. Rong, and J. F. Du (2019); A. Pick, S. Silberstein, N. Moiseyev, and N. Bar-Gill (2019). Quantum effects including photon blockadeR. Huang, S. K. Özdemir, J. Q. Liao, F. Minganti, L. M. Kuang, F. Nori, and H. Jing (2022); Z. G. Li, X. M. Li, G. F. Zhang, and X. L. Zhong (2023); Y. L. Zuo, R. Huang, L. M. Kuang, X. W. Xu, and H. Jing (2022); Z.-H. Yuan, Y.-J. Chen, J.-X. Han, J.-L. Wu, W.-Q. Li, Y. Xia, Y.-Y. Jiang, and J. Song (2023) and topological quantum state controlW. Q. Liu, Y. Wu, C. K. Duan, X. Rong, and J. F. Du (2021); Z. Gong, Y. Ashida, K. Kawabata, K. Takasan, S. Higashikawa, and M. Ueda (2018); H. Xu, D. Mason, L. Y. Jiang, and J. G. E. Harris (2016) have been achieved by engineering the EPs. As a most important quantum property, EP-induced exceptional entanglement behaviors have also been observed in very recent works. For example, the occurrence of the entanglement transition at the EP has been observed in a NH system composed of qubit coupled with photonset al. (2023). Entanglement generation between two NH qubits can be accelerated by approaching the EP of the systemZ.-Z. Li, W. Chen, M. Abbasi, K. W. Murch, and K. B. Whaley (2023). Replacing one NH qubit with unitary qubit, entanglement can be maximized at the EPA. Kumar, K. W. Murch, and Y. N. Joglekar (2022). However, the relationship between the EPs and entanglement dynamics remains unknown.
@@
Here, we show that multitype entanglement dynamics can be realized by engineering the EPs, by analyzing a generic model composed of two coupled NH qubits. The inter-qubit coupling will lower the order of EP from the original fourth-order EP (EP4) to a second-order EP (EP2). However, we find that disparate entanglement dynamics on the two sides of the original EP4 and EP2 can be observed simultaneously in the case of weak coupling (i.e., the coupling strength is much smaller than the dissipation rate), indicating different types of entanglement behaviors can be induced by the EPs. When enhancing the coupling strength to the strong coupling regime (i.e., the coupling strength and dissipation rate are on the same order), the original EP4 is replaced by an additional EP2, leading to the disappearance of the EP4-engineered entanglement changing. Taking Ising type interaction as an example, we also find EP-induced entanglement changing in the absence of the driving field. Our scheme has good experimental feasibility, different types of platforms can be used to study the simulation of non-Hermitian systems and the observation of EPs, such as the coupled loss and gain cavities Hodaei et al. (2014); Smirnov et al. (2018), cavity optomechanical systems Zhang et al. (2018); Wu et al. (2023), superconducting circuits M. Naghiloo, M. Abbasi, Y. N. Joglekar, and K. W. Murch (2019); A. Blais, S. M. Girvin, and W. D. Oliver (2020); Z. L. Xiang, S. Ashhab, J. Q. You, and F. Nori (2013); M. H. Devoret and R. J. Schoelkopf (2013), ion traps C. Monroe and J. Kim (2013); et al. (2021) and single-spin system Y. Wu, W. Q. Liu, J. P. Geng, X. R. Song, X. Y. Ye, C. K. Duan, X. Rong, and J. F. Du (2019). We believe that our study provides an efficient method for designing EP based quantum devices, such as the quantum-state synthesis using exceptional-point engineering in superconducting resonators W. S. Teixeira, V. Vadimov, T. Mörstedt, S. Kundu, and M. Mottonen (2023), single-photon interferometric network et al. (2017); L. Xiao, K. K. Wang, X. Zhan, Z. H. Bian, K. Kawabata, M. Ueda, W. Yi, and P. Xue (2019) and enhancing the performance of quantum metrology near the EP et al. (2016).
The paper is organized as follows. Section II presents theoretical models for the two coupled non-Hermitian qubits. In Section III, we conduct numerical simulations for predicting the EPs in the system. We show the entanglement dynamics change in Section IV and discuss the reason of this phenomenon in Section V. Furthermore, we demonstrate that the entanglement dynamics persist in systems with dipolar coupling, as discussed in Section VI and VII, respectively. We summarize our findings in Section VIII.
@@
Figure 1: Schematic diagram of the multitype entanglement dynamics (I, II, III) induced by EPs with increasing $\Omega$. The regions of PTB and PTS represent the $\mathcal{P}\mathcal{T}$-symmetry-broken phase and the $\mathcal{P}\mathcal{T}$-symmetric phase, respectively. The EPs can be flexibly changed by tuning the drive amplitude $\Omega$. (a) The system exhibits a fourth-order EP (EP4) without coupling and no entanglement will be generated in this case. (b) In the weak coupling regime (the coupling strength $\xi$ is much smaller than the dissipation rate $\gamma$ of the qubits, i.e., $\xi\ll\gamma$), the system exhibits the original EP4 (EP4’) and an existing second-order EP (EP2). (c) In the strong coupling regime (i.e, $\xi\sim\gamma$), the original EP4 disappears and an additional EP2 emerges. Corresponding to the presence of EP4’, the purely imaginary (PI) eigenvalues become a mixture of both real and imaginary parts. The EP2 induces the spectrum transition of the PTS and the mixed phase. The orange, white and blue represent the region of the PI, PTS and mixed phase, respectively. The entanglement behavior of type I in PI is monotonically increasing and eventually reaching a stable value, type II in PTS corresponds to the entanglement behavior of continuous oscillations over the entire time domain, and type III in the mixed phase combines the characteristics of the above two different entanglement behaviors.
@@ II Model
${{H}_{j=1,2}}=\sum\limits_{j=1,2}{\left(-\frac{{{\gamma}_{j}}}{2}\sigma_{j}^{-}\sigma_{j}^{+}+{{\Omega}_{j}}\sigma_{j}^{x}\right)+{{H}_{\operatorname{int}}}}.$ (1)
@@
The Pauli operators in terms of the quantum energy levels-with no classical analogs- ${{\left|b\right\rangle}_{j}}$ and ${{\left|a\right\rangle}_{j}}$ as $\sigma_{j}^{+}={{\left|a\right\rangle}_{j}}\left\langle b\right|,\sigma_{j}^{-}={{\left|b\right\rangle}_{j}}\left\langle a\right|$ and $\sigma_{j}^{x}={{\left|b\right\rangle}_{j}}\left\langle a\right|+{{\left|a\right\rangle}_{j}}\left\langle b\right|,j\in\{1,2\}$. ${{\gamma}_{j}}$ and ${{\Omega}_{j}}$ are the energy decay rate of ${{\left|b\right\rangle}_{j}}$ and driving amplitude, respectively. In the following analysis, we assume resonance between the driving field frequency and the qubit transition, with ${{\gamma}_{1}}={{\gamma}_{2}}=\gamma$ and ${{\Omega}_{1}}={{\Omega}_{2}}=\Omega$ for simplicity. The third term represents the interaction between the two NH qubits. When the two qubits are coupled via dipolar interaction, the Hamiltonian can be described as ${{H}_{\operatorname{int}}}=g\left(\sigma_{1}^{\dagger}{{\sigma}_{2}}+{{\sigma}_{1}}\sigma_{2}^{\dagger}\right)$, which can be realized in various systems such as superconducting circuit systems M. Naghiloo, M. Abbasi, Y. N. Joglekar, and K. W. Murch (2019); M. Dalmonte, S. I. Mirzaei, P. R. Muppalla, D. Marcos, P. Zoller, and G. Kirchmair (2015) and Rydberg atom systems S. de Léséleuc, V. Lienhard, P. Scholl, D. Barredo, S. Weber, N. Lang, H. P. Büchler, T. Lahaye, and A. Browaeys (2019). In addition to this interaction type, there is another type of interaction between the two qubits or spins, i.e., Ising type interaction, which can be represented as ${{H}_{\operatorname{int}}}=-\xi{{\left(\sigma_{1}^{x}+\sigma_{2}^{x}\right)}^{2}}$. Here $g$ and $\xi$ denote the coupling strength, respectively. Ising type interaction has been widely studied in hybrid spin-mechanical system X.-F. Pan, X.-L. Hei, X.-L. Dong, J.-Q. Chen, C.-P. Shen, H. Ali, and P.-B. Li (2023); P. B. Li, Z.-L. Xiang, P. Rabl, and F. Nori (2016), the quantum spin lattice systems W. Dür, L. Hartmann, M. Hein, M. Lewenstein, and H. J. Briegel (2005); J. W. Britton, B. C. Sawyer, A. C. Keith, C. C. J. Wang, J. K. Freericks, H. Uys, M. J. Biercuk, and J. J. Bollinger (2012) and hybrid circuit systemsJ. G. Bohnet, B. C. Sawyer, J. W. Britton, M. L. Wall, A. M. Rey, M. Foss-Feig, and J. J. Bollinger (2016); M. Block, Y. Bao, S. Choi, E. Altman, and N. Y. Yao (2022). In this work, we focus on the entanglement dynamics change induced by the EPs, and the main results are summarized in Fig. 1.
III Energy spectrum and EPs
@@
In the following we mainly focus on the case of Ising type interaction (the calculation of the dipolar interaction case can be found in Sec. VI). The system Hamiltonian in Eq. (1) reduces to a so-called passive $\mathcal{P}\mathcal{T}$-symmetric Hamiltonian A. Guo, G. J. Salamo, D. Duchesne, R. Morandotti, M. Volatier-Ravat, V. Aimez, G. A. Siviloglou, and D. N. Christodoulides (2009) can be written as $H={{H}_{\mathcal{PT}}}-\left(i\gamma/2\right)I$ with $I$ being the unitary matrix. The Hamiltonian ${{H}_{\mathcal{PT}}}$ can be written as
@@
$\begin{split}{{H}_{\mathcal{PT}}}=&\left(\begin{matrix}-2\xi+\frac{i\gamma}{2}&\Omega&\Omega&-2\xi\\ \Omega&-2\xi&-2\xi&\Omega\\ \Omega&-2\xi&-2\xi&\Omega\\ -2\xi&\Omega&\Omega&-2\xi-\frac{i\gamma}{2}\\ \end{matrix}\right).\end{split}$ (2)
This Hamiltonian satisfies $\mathcal{P}\mathcal{T}$ symmetry, where the parity operator is given as $\mathcal{P}=\sigma_{1}^{x}\sigma_{2}^{x}$. The eigenvalues of the system without coupling, i.e., $\xi=0$ ($g=0$), are ${{\lambda}_{0,1}}=-i\gamma/2$ and ${{\lambda}_{\pm}}=-i\gamma/2\pm i\sqrt{16{{\Omega}^{2}}-{{\gamma}^{2}}}/2$, indicating that the system has a four-order EP (EP4) if $\Omega={{\Omega}_{\text{EP4}}}=\gamma/4$ [Figs. 2(a) and  2(d)]. In the region of $\mathcal{P}\mathcal{T}$-symmetric phase (PTS) (i.e., $\Omega>{{\Omega}_{\text{EP4}}}$), the eigenvalues of the system are purely real, while the eigenvalues are purely imaginary (PI) in the region of $\Omega<{{\Omega}_{\text{EP4}}}$, this phase region is defined as PI [see Fig. 2(a, d)]. As described in Fig. 2(b) and  2(e), an EP2 can be found at ${{\Omega}_{\text{EP2}}}/\gamma=0.283$ when choosing $\xi/\gamma=0.006$ as an example. Although the EP4 disappears in the weak coupling regime (i.e., $\xi\ll\gamma$), it still has influence on the quantum effect. Thus, the original EP4 (i.e., EP4’) can be read as the “ghost” of the factual EP4. Corresponding to the presence of EP4’, the purely imaginary (PI) eigenvalues become a mixture of both real and imaginary parts. The EP2 induces the spectrum transition of the PTS and the mixed phase [see the blue regions in Fig. 2(b, e)]. It is well-known that in the region of $\mathcal{P}\mathcal{T}$-symmetry-broken phase (PTB), the eigenvalues are a pair of conjugate complex numbers, that is $\lambda=-{{\lambda}^{*}}$. Therefore, both the mixed phase and PI belong to the PTB.
@@
The order of EP will be lowered from 4 to 2 when adding the qubit-qubit coupling. An additional EP2 will appear in the system when enhancing the coupling strength to the strong coupling regime (i.e., $\xi\sim\gamma$). As demonstrated in Fig. 2(c) and  2(f), one EP2 appears at the critical value $\Omega_{\text{EP2}}^{(1)}/\gamma=0.626$ and an additional EP2 is at $\Omega_{\text{EP2}}^{(2)}/\gamma=1.34$ when choosing $\xi/\gamma=0.5$. The region of the mixed phase occurs at $\Omega_{\text{EP2}}^{(1)}<\Omega<\Omega_{\text{EP2}}^{(2)}$ and the regions of PTS take place at $\Omega<\Omega_{\text{EP2}}^{(1)}$ and $\Omega>\Omega_{\text{EP2}}^{(2)}$, which indicates that the mixed phase is surrounded by the two EP2s in the strong coupling regime.
@@
Figure 2: Real (a, b, c) and imaginary (d, e, f) parts of eigenvalues for different Ising type interaction strength. (a) Real and (d) imaginary parts of the eigenvalues for $\xi/\gamma=0$. As examples for weak coupling regime for $\xi/\gamma=0.006$ (b, e) and strong coupling regime for $\xi/\gamma=0.5$ (c, f), respectively. The two qubits are assumed that have the same drive amplitude $\Omega$ and the same decay rates $\gamma$. The orange and white regions represent the PI and PTS, respectively. The mixed phase is formed by the original EP4 and existing EP2 (i.e., ${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$) in the case of weak coupling, while is surrounded by the two EP2s (i.e., $\Omega_{\text{EP2}}^{(1)}<\Omega<\Omega_{\text{EP2}}^{(2)}$) in the strong coupling regime, as illustrated in the blue regions.